OpenStudy (anonymous):

Solve: negative 1 over 2 x + 1 = –x + 8 7 1 over 2 14 18

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

\[-\frac{1}{2}x+1=-x+8\] like that?

OpenStudy (dr0zier99):

x=14

OpenStudy (anonymous):

don't give out a direct answer because then they wont be learning how to solve the problem that they cant solve :) @ginapondd ?

OpenStudy (misty1212):

lol you can just pick C each time and get a score of about 82%